# The Cognitive Foundations of Economic Exchange: A Modular Framework Grounded in Behavioral Evidence

### v2

@@ Abstract
A key challenge in multi-agent AI is modeling social cooperation under realistic behavioral constraints. Many foundational concepts in economics and ethics—such as ”trust” or ”morality”—are often defined informally, without operational criteria or cognitive grounding, which limits their testability and implementation in artificial agents. Drawing on converging empirical evidence from primate behavior, infant cognition, and economic anthropology, we propose a conceptual framework composed of three cognitively minimal mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, enabling the bottom-up emergence of scalable cooperation, proto-economic norms, and institutional structure.
@@ 1 Introduction
A persistent gap in economic simulation is the lack of behavioral grounding: most models begin with institutions—markets, contracts, payoff matrices—while assuming the cognitive and social mechanisms that make such systems viable. This limits both the explanatory power and simulateability of artificial economic agents.
This issue is mirrored in broader AI+Society research. Concepts like “trust,” “value,” and “cooperation” are often treated as abstract labels, operationalized only at the outcome level. Without grounding these constructs in observable mechanisms, models risk simulating the surface of economic life without modeling its origins.
In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, behaviorally grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. These capacities are observed across human and nonhuman primates, and support sustained reciprocal exchange without requiring symbolic trust or formal enforcement.
Anthropological accounts have long shown that early human exchange was not rooted in barter or equivalence, but in socially embedded reciprocity. Systems like the Kula ring (Malinowski, 2013) emphasized obligation and alliance over transaction. Similarly, primatology has documented robust reciprocal behaviors among chimpanzees and bonobos—grooming (De Waal, 1997), food sharing, and coalition—indicating that the behavioral foundations of reciprocal exchange are observable across species and precede formal market systems.
We propose a modular framework built from these three mechanisms, designed to serve as behavioral primitives in economic simulation. This approach not only bridges cognitive science and institutional theory, but also enables bottom-up modeling of how debt, credit, and monetary behavior may emerge from repeated social interaction.
@@
We offer a theoretical realignment: a cognitively grounded account of reciprocal exchange that clarifies how economic systems can emerge from social behavior rather than institutional design. We argue that reciprocity is not just a component of cooperation, but the minimal cognitive substrate from which stable social structures emerge.
@@
  • Reinterpret key findings from behavioral economics—such as inequity aversion, reciprocal cooperation, and altruistic punishment—as consistent expressions of underlying biological mechanisms, rather than anomalies to rational choice theory.
Unlike prior models that treat cooperation heuristically or assume trust as exogenous, our framework offers a cognitively minimal and simulateable account of how such expectations emerge from social interaction. This work reframes economic modeling not as a top-down implementation of rules, but as a bottom-up emergence of institutions from cognitive interaction.
@@ 2.1 Agent-Based Social Simulation
Recent advances in multi-agent language models have produced systems capable of planning, negotiation, and scripted cooperation (Park et al., 2023; Li et al., 2023). However, these models typically lack behavioral grounding: they simulate interaction as token-level dialogue without modeling the internal mechanisms that make social exchange stable—such as partner memory, reciprocity tracking, or cost–benefit sensitivity.
While some frameworks include persistent memory modules or heuristic-based reasoning, these components often lack principled constraints. Agents either memorize indiscriminately or update state in ways that are unstable or opaque, making it difficult to simulate long-term reciprocal dynamics. As a result, apparent cooperation often emerges from prompt bias or hardcoded behavior, not from simulateable social inference.
This gap has two key consequences: (1) It limits the explanatory value of multi-agent simulations as models of human social behavior, and (2) it prevents systematic exploration of how early exchange mechanisms scale into institutional forms. Our framework addresses this by identifying minimal, biologically plausible mechanisms that can be embedded in LLM agents to support grounded, dynamic reciprocity over time.
@@ 2.2 Reciprocal Behavior in Nonhuman Primates
Research in primatology provides critical insight into the behavioral foundations of reciprocity. Species such as chimpanzees and bonobos engage in reciprocal behaviors across contexts like food sharing, grooming, and coalition support. For instance, de Waal (De Waal, 1997) illustrated how chimpanzees form long-term social alliances maintained through reciprocal exchange, and Brosnan et al. (Brosnan & De Waal, 2003) demonstrated that primates are sensitive to fairness and inequity in outcomes.
While these studies have richly documented the behavioral forms of reciprocity, they often stop short of linking such behaviors to the broader architecture of economic interaction. The question of how these mechanisms scale or evolve into structured systems of exchange remains relatively underexplored.
Our framework builds on these findings by situating primate reciprocity within a cognitive-behavioral model. We argue that capacities such as individual recognition, memory of interaction history, and cost-return sensitivity form a behaviorally grounded substrate upon which more complex human economic systems could emerge.
@@ 2.3 Social Exchange Without Markets
Anthropological accounts have long challenged the barter-origin myth that underpins classical economic theory. Foundational works such as The Gift (Mauss, 2024) and Stone Age Economics (Sahlins, 2013) argue that early exchange was not driven by market logic, but embedded in webs of social relationships, mutual obligation, and symbolic prestige.
These accounts have been instrumental in shifting attention away from transactional equivalence and toward socially embedded forms of value. Yet, they often bracket the question of what kinds of cognitive and behavioral mechanisms make such systems feasible. In other words: what allows individuals to engage in sustained, deferred, and socially contingent exchange?
Our framework complements these anthropological insights by identifying the minimal behavioral architecture that enables reciprocity to stabilize across time and social partners. We argue that reciprocity is not merely a cultural invention, but a behaviorally tractable structure observable across species.
2.4 From Behavioral Anomalies to Cognitive Substrates
Behavioral economics has significantly advanced our understanding of social preferences, particularly in contexts of fairness, cooperation, and reciprocity. Paradigms such as the Ultimatum Game (Güth et al., 1982), Trust Game (Berg et al., 1995), and Public Goods Game (Fehr & Gächter, 2002) reveal systematic deviations from self-interested rationality, with participants rejecting unfair offers, rewarding cooperation, and punishing free-riders.
These empirical patterns have led to influential models of “inequity aversion” (Fehr & Schmidt, 1999) and “strong reciprocity” (Gintis, 2000), which posit internalized fairness norms and preferences for norm enforcement. However, these models typically treat the relevant capacities—such as fairness sensitivity or social tracking—as given, rather than as targets of explanation.
In contrast, our framework treats these preferences as emergent outcomes of specific cognitive-social mechanisms: individual recognition, memory of interaction, and cost–return evaluation. By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a behaviorally grounded account of exchange behavior.
@@ 3.1 The Myth of Barter and the Reciprocal Foundations of Exchange
The textbook account of economic origins often begins with barter: the notion that early humans exchanged goods directly in the absence of currency, with markets and money arising later to reduce transactional friction. This narrative, while intuitively appealing, has little support in either ethnographic or historical records.
As Sahlins argued in Stone Age Economics (Sahlins, 2013), early exchange systems were not structured around equivalence, but around relational modes of reciprocity—generalized, balanced, and negative. These forms differ not in rationality but in social distance and temporal orientation: generalized reciprocity, typical among kin, involved open-ended giving without expectation of immediate return; balanced reciprocity, among peers, entailed roughly equivalent exchanges over time; and negative reciprocity, often among rivals or strangers, reflected opportunistic or exploitative behavior.
A paradigmatic ethnographic example is the Kula ring, described by Malinowski (Malinowski, 2013), in which ceremonial armbands and necklaces circulate across Trobriand Islands through long-term, prestige-based exchanges. These objects held little utilitarian value, yet were exchanged with strict directional patterns, social memory, and partner exclusivity. What appears as a symbolic ritual is, upon closer inspection, a complex system of relational tracking, reputational assessment, and obligation extension—precisely the kinds of cognitive demands that structure early forms of exchange.
Mauss’s The Gift (Mauss, 2024) further formalized this insight by showing how gift-giving in archaic societies is embedded in a triadic obligation: to give, to receive, and to reciprocate. The gift is not merely a transfer of value, but a means of sustaining long-term social bonds. When viewed from this angle, early exchange is not the product of impersonal rational choice, but of biologically and socially grounded behaviors—tracking individuals, remembering interactions, maintaining alliances, and adjusting for asymmetrical returns.
These foundations suggest that exchange predates markets not just historically, but cognitively. What emerges from these systems is not merely institutional scaffolding, but a structured behavioral logic—one that we formalize in the following section.
@@ 3.2 The Three Behavioral Mechanisms Enabling Reciprocity
If the earliest forms of economic interaction were not driven by impersonal barter, then what behavioral dynamics made sustained exchange possible? We argue that reciprocal exchange emerged not from a need to equalize value, but from the ability to maintain accountability and continuity across repeated social interactions.
These early exchanges were not market transactions, but embedded social patterns shaped by memory, mutual responsiveness, and partner-specific interaction history. Exchange, in this view, arises not from price formation, but from behavioral infrastructure: the cognitive and social capacities that allow individuals to track others, assess cooperative value, and anticipate reciprocity.
To formalize these behavioral foundations, we propose a cognitively grounded framework composed of three simulateable mechanisms:
  • Individual recognition: enabling agents to identify and re-engage specific social partners over time;
  • Reciprocal credence: a scalar, updateable expectation that cooperation will be returned;
  • Cost–return sensitivity: allowing agents to regulate cooperative behavior based on dynamic payoff asymmetries.
These mechanisms are supported by converging evidence from nonhuman primates and human infants, suggesting they are biologically plausible, developmentally early, and cognitively minimal. Each is also behaviorally tractable and simulateable in multi-agent systems—whether as memory modules, value estimators, or policy regulators.
While primate studies have richly documented behaviors like grooming, food sharing, and coalition support, the cognitive prerequisites for sustaining these interactions are often left implicit. As (Schino & Aureli, 2010) note, reciprocity is frequently treated as a local response pattern, rather than a structured process involving memory and expectation. Yet reciprocal exchange, by definition, requires continuity. Without recognition or anticipation of future interaction, trust-like behavior cannot functionally emerge.
By defining these three mechanisms as simulateable behavioral primitives, we provide a bottom-up framework for modeling the emergence of economic exchange—one that avoids presupposing institutions, markets, or symbolic trust, and instead builds from capacities that are both empirically grounded and computationally implementable.
Figure 3: The agent integrates three input pathways via behaviorally grounded mechanisms—recognition, credence, and cost–return evaluation—to support dynamic decisions in social exchange contexts.
@@ 4.1 Individual Recognition
Individual recognition is the capacity to identify and distinguish specific social partners across time. It enables agents to track prior interactions with particular individuals, forming the basis for memory-based expectations about future behavior.
This capacity underlies the first and most basic source of reciprocal credence: without individual recognition, no information about past cooperation can be meaningfully retained or applied to guide future interaction.
@@
Among nonhuman primates, this capacity is especially robust. A recent study shows that chimpanzees and bonobos remember familiar individuals after decades of separation, and preferentially attend to those with whom they had positive social histories (Lewis et al., 2023). These findings show that individual recognition in apes is long-term and identity-specific, enabling the tracking of social partners across time.
In nonhuman primates, individual recognition enables partner-specific cooperation. In rope-pulling tasks, chimpanzees adjusted their actions based on the identity and prior behavior of their partners—pausing, glancing, or gesturing to initiate coordination (Hirata & Fuwa, 2007). Dyads with higher food-sharing tolerance were also more successful in joint tasks, indicating that cooperative memory is structured around known social partners(Melis et al., 2006b).
In humans, this ability emerges early. Infants as young as 14 months selectively help or share with individuals who have previously acted prosocially (Dunfield et al., 2011). Such findings suggest that recognition, even in its early form, is tied to social history and used to guide future cooperative behavior.
@@ 4.2 Reciprocal Credence
The term “trust” is widely used across disciplines, yet rarely defined with precision. In both everyday language and scientific literature, it is invoked to describe moral commitment, emotional closeness, institutional reliability, and behavioral expectation—often interchangeably. This conceptual ambiguity makes trust difficult to operationalize, especially in comparative, developmental, or artificial settings.
To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded belief that another agent will respond to cooperation with cooperation. Unlike categorical trust, reciprocal credence is scalar, updateable, and context-sensitive. It is not a feeling, but a functional belief state—one that governs an agent’s willingness to initiate or sustain cooperative behavior.
@@
Reciprocal credence is a scalar, updateable belief that another agent will respond to cooperation with cooperation. It is not a feeling, but a functional decision variable that governs an agent’s willingness to initiate or sustain cooperative behavior.
@@
  • Prospective inference, based on contextual signals that indicate future reciprocity is likely—such as repeated interaction, social roles, or institutional incentives.
@@
    Reputational inference, based on observed or reported behavior toward others—supporting third-party inference of prosocial disposition.
@@
These pathways differ in reliability, cognitive cost, and developmental trajectory—but together, they provide the informational grounding for reciprocal credence as a simulateable and biologically plausible substrate for reciprocal cooperation.
@@
Empirical studies support the differential emergence of these sources across species and developmental stages. In chimpanzees, reciprocal cooperation is consistently observed when individuals have direct histories of prosocial interaction, with little evidence for reputation-based inference or role-based expectation. Similarly, human infants display early sensitivity to agents who have helped them directly, yet show limited use of third-party observation or role generalization. These patterns suggest that direct positive interaction is the foundational layer of reciprocal credence, while the other two sources emerge developmentally and culturally.
Across species, direct prosocial interaction serves as the most robust and foundational basis for reciprocal credence. In wild chimpanzees, grooming is exchanged not merely in-the-moment, but over extended time horizons. Longitudinal data reveal that grooming given and received is significantly correlated across dyads, even with delays spanning several days—indicating that chimpanzees maintain memory of past interactions and calibrate future behavior accordingly (Gomes et al., 2009). Experimental studies further show that chimpanzees prefer to collaborate with previous cooperative partners, suggesting they encode individual-specific social history and deploy it strategically in cooperative contexts (Melis et al., 2006a).
Human infants exhibit a comparable early capacity. By three years of age, children are less likely to help individuals who have previously acted with harmful intent—even when the harm was directed at a third party, and even if no harm was successfully done. This selective avoidance indicates that infants track not only outcomes but also the moral valence of agents’ prior actions when deciding to engage prosocially (Vaish et al., 2010).
Together, these findings support the primacy of direct positive interaction in generating reciprocal credence. Unlike inference or abstraction, this source relies on lived experience—anchoring cooperative expectations in observable history, not hypothetical roles or indirect cues.

### v3

@@ Abstract
The origins of economic behavior remain unresolved—not only in the social sciences but also in AI, where dominant theories often rely on predefined incentives or institutional assumptions. Contrary to the longstanding myth of barter as the foundation of exchange, converging evidence from early human societies suggests that reciprocity—not barter—was the foundational economic logic, enabling communities to sustain exchange and social cohesion long before formal markets emerged. Yet despite its centrality, reciprocity lacks a simulateable and cognitively grounded account. Here, we introduce a minimal behavioral framework based on three empirically supported cognitive primitives—individual recognition, reciprocal credence, and cost–return sensitivity—that enable agents to participate in and sustain reciprocal exchange, laying the foundation for scalable economic behavior. These mechanisms scaffold the emergence of cooperation, proto-economic exchange, and institutional structure from the bottom up. By bridging insights from primatology, developmental psychology, and economic anthropology, this framework offers a unified substrate for modeling trust, coordination, and economic behavior in both human and artificial systems.
@@ 1 Introduction
The standard origin story of economics begins with barter: individuals trading goods directly, with money and markets emerging to reduce friction. Yet this narrative is a myth. Ethnographic and historical research finds little evidence for barter as a primary mode of early exchange (Sahlins, 2013; Mauss, 2024; Malinowski, 2013). Instead, early human societies were structured around reciprocity—a temporally extended system of giving, receiving, and returning, embedded in social relationships.
Surprisingly, most economic models still begin where barter leaves off: with institutions, contracts, and payoff matrices. These frameworks assume—but do not explain—the behavioral substrate that makes cooperation possible in the first place. As a result, we lack simulateable accounts of how economic behavior emerges from interaction among agents.
This blind spot extends to broader AI + society research. Core concepts like “trust,” “value,” and “cooperation” are often used without precise definitions or behavioral grounding. Rather than being modeled as well-defined, simulateable mechanisms, they are treated as vague abstractions—obscuring rather than explaining the dynamics of social interaction.
In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—emerge not from barter or institutional design, but from reciprocity: a structured pattern of interaction built on three cognitively minimal mechanisms— individual recognition, reciprocal credence, and cost–return sensitivity. These mechanisms, observed across humans and nonhuman primates, support partner-specific cooperation without relying on symbolic trust, contracts, or centralized enforcement.
We formalize these mechanisms as simulateable behavioral primitives, providing a bottom-up framework grounded in cognitive and behavioral evidence. By integrating insights from anthropology, developmental psychology, and economic anthropology, our approach offers a biologically grounded alternative to institution-first models of cooperation.
@@
We provide a minimal theoretical framework grounded in behaviorally observable mechanisms. This framework yields four core contributions:
@@
  • Reinterpret classic behavioral findings—such as framing effects, loss aversion, and anchoring bias—as consistent expressions of bounded, biologically grounded cognition, rather than anomalies to rational choice theory.
Unlike models that treat trust heuristically or assume cooperation as exogenous, our framework offers a minimal, simulateable foundation for how cooperative expectations and exchange structures emerge from interaction. This allows us to reframe the origins of economic systems not as top-down implementations of institutions, but as bottom-up constructions grounded in cognitive behavior.
@@ 2.1 Agent-Based Social Simulation
Recent work on multi-agent language models has enabled scripted cooperation and task planning (Park et al., 2023; Li et al., 2023), but these systems rarely model the cognitive mechanisms that support stable social interaction—such as partner tracking, interaction memory, or sensitivity to prior outcomes.
While some frameworks add memory or heuristics, these are typically unconstrained and ad hoc, limiting their ability to simulate reciprocal dynamics over time. Apparent cooperation often stems from prompt bias or hardcoded behavior rather than simulateable social inference.
Our framework addresses this gap by identifying minimal, biologically grounded mechanisms that support dynamic reciprocity, enabling more realistic modeling of social behavior and scalable exchange.
@@ 2.2 Reciprocal Behavior in Nonhuman Primates
Studies in primatology reveal that species such as chimpanzees and bonobos engage in reciprocal acts across grooming, food sharing, and alliance formation. For instance, de Waal (De Waal, 1997) documented long-term cooperative relationships maintained through delayed reciprocity, while Brosnan et al. (Brosnan & De Waal, 2003) showed sensitivity to fairness and outcome inequity.
Although these findings offer strong evidence of reciprocal behavior, they are seldom linked to broader economic structures or coordination dynamics. The underlying cognitive mechanisms—and how they scale to support complex social systems—remain largely unexplored.
@@ 2.3 Social Exchange Without Markets
Anthropological work has challenged the classical view that early economies began with barter. Foundational accounts, such as Mauss’s The Gift (Mauss, 2024) and Sahlins’s Stone Age Economics (Sahlins, 2013), emphasize that early exchange was embedded in networks of kinship, obligation, and prestige—rather than driven by market logic or equivalent trade.
These studies highlight the importance of social context and long-term reciprocity, but often leave open the question of what mechanisms make such systems sustainable. Specifically, the cognitive conditions under which deferred, partner-contingent exchange can stabilize across time are rarely formalized.
@@ 3.1 The Myth of Barter and the Reciprocal Foundations of Exchange
The standard narrative of economic origins begins with barter: the idea that early humans exchanged goods directly, with markets and money emerging later to reduce friction. But this account lacks support in ethnographic and historical records.
As Sahlins argued in Stone Age Economics (Sahlins, 2013), early exchange was not based on equivalence but on relational reciprocity: generalized reciprocity involved open-ended giving among kin; balanced reciprocity entailed delayed equivalence among peers; and negative reciprocity reflected opportunism among strangers. Malinowski’s Kula ring (Malinowski, 2013) exemplifies long-term, prestige-based exchange using non-utilitarian objects, governed by memory, exclusivity, and directional flow. Beneath the ritual lies a cognitively rich structure of tracking, obligation, and reputation. Mauss’s The Gift (Mauss, 2024) formalized this logic as a triadic obligation—to give, receive, and reciprocate—framing gift exchange as a cognitive and social mechanism for sustaining long-term bonds.
Together, these accounts suggest that early human societies were not built on barter, but on reciprocity: a temporally extended structure of interaction grounded in memory, obligation, and repeated engagement.
Far from a peripheral exchange strategy, reciprocity was the behavioral infrastructure of early economies—the cognitive and social foundation that enabled resource flow, social cohesion, and relational continuity.
It sustained cooperation long before the emergence of contracts or currency, and predates markets not only historically, but cognitively—anchored in the minimal behavioral logic we formalize in the next section.
@@ 3.2 The Three Behavioral Mechanisms Enabling Reciprocity
If reciprocity—not barter—underpins early economic life, its roots must predate formal institutions. This raises a fundamental question: what minimal behavioral and cognitive mechanisms are sufficient to sustain reciprocal exchange?
Nonhuman primates offer a critical comparative lens. In chimpanzees and bonobos, sustained, partner-contingent behaviors—such as food sharing, grooming, and coalition support—exhibit structured patterns of reciprocity and serve as core mechanisms for maintaining group cohesion and social stability (De Waal, 1997).
Comparative psychology isolates the minimal cognitive substrates that support such behaviors. By examining species with simpler interaction structures, it avoids cultural and institutional confounds—revealing the baseline capacities sufficient for reciprocity to emerge and stabilize.
To formalize these foundations, we introduce a cognitively grounded framework comprising three simulateable mechanisms:
  • Individual recognition: identifying and re-engaging specific social partners over time;
  • Reciprocal credence: an updateable expectation that cooperation will be returned;
  • Cost–return sensitivity: modulating cooperative behavior in response to payoff asymmetries.
We treat these as simulateable behavioral primitives—providing a bottom-up foundation for modeling the emergence of economic exchange, without presupposing institutions, markets, or symbolic trust.
Figure 3: The agent integrates three input pathways via behaviorally grounded mechanisms—recognition, credence, and cost–return evaluation—to support dynamic decisions in reciprocal exchange.
@@ 4.1 Individual Recognition
Individual recognition is the capacity to identify specific social partners over time, enabling agents to track past interactions and form expectations about future behavior. It is the prerequisite for memory-based reciprocity: without it, no past cooperation can inform future decisions.
@@
Nonhuman primates show strong individual recognition. Chimpanzees and bonobos remember familiar individuals after decades and preferentially attend to former social partners (Lewis et al., 2023).
In cooperative tasks, chimpanzees adjust behavior based on partner identity and prior interactions—initiating coordination through glances, pauses, or approach behaviors—and succeed more often with tolerant, familiar partners (Hirata & Fuwa, 2007; Melis et al., 2006b).
In humans, this ability emerges in infancy. By 14 months, children selectively help those who previously acted prosocially (Dunfield et al., 2011), suggesting that recognition is tied to social memory and used to guide future cooperation.
@@ 4.2 Reciprocal Credence
The term “trust” is widely used across disciplines, yet rarely defined with precision. It is invoked to describe moral commitment, emotional closeness, institutional reliability, and behavioral expectation—often interchangeably. This ambiguity makes it difficult to operationalize trust, especially in comparative, developmental, or artificial settings.
To address this, we introduce the concept of reciprocal credence:
@@
Reciprocal credence is a graded, updateable belief that cooperative behavior will be reciprocated. It functions as a decision variable governing an agent’s willingness to initiate or sustain cooperation.
@@
  • Prospective inference, based on contextual signals that indicate future reciprocity is likely—such as social roles or reputational cues.
@@
    Third-party inference, based on observed or reported behavior toward others—supporting third-party inference of prosocial disposition.
@@
Together, these sources form a minimal, biologically grounded substrate for modeling reciprocal behavior in both human and artificial agents.
@@
Direct positive interaction provides the most robust foundation for reciprocal credence. In chimpanzees, grooming is exchanged not only immediately but across time: dyads reciprocate over multi-day delays, suggesting memory-based calibration of cooperation (Gomes et al., 2009). They also preferentially collaborate with partners who have helped them before (Melis et al., 2006a), indicating an ability to track and evaluate individual-specific cooperative history.
Similarly, by age three, human children avoid helping agents who previously harmed others, even if the harm was unsuccessful (Vaish et al., 2010)—suggesting early sensitivity to past intent in guiding prosocial choice.
By contrast, third-party inference and role-based expectation are rarely observed in nonhuman primates and emerge only later in human development. Nonetheless, both are deeply embedded in human social life, forming the basis for reputation systems, institutional roles, and generalized cooperation beyond direct experience.